how much of an alloy that is 40 copper should be mixed with 100 ounces of an alloy that is 90% copper in order to get an alloy that is 50% copper?
Answer by lwsshak3(11628)

how much of an alloy that is 40 copper should be mixed with 100 ounces of an alloy that is 90% copper in order to get an alloy that is 50% copper?
***
let x=amount 0f 40% copper to be mixed
x+100=amt of 50% alloy mixture
40%x+90%*100=50%(x+100)
.4x+90=.5x+50
.10x=40
x=400
amount of 40% copper to be mixed with 100 ounces of an alloy that is 90% copper=400 ounces